# Revisiting Static Feature-Based Android Malware Detection

### v1

@@ 4.1.3 Duplicate Statistics in the Datasets
In the validation and test sets, we observe two key patterns regarding duplicates. First, the fraction of duplicate samples varies monthly. Second, malware samples tend to contain more duplicates than benign samples, with certain months particularly affected. For instance, in the first test month of the Drebin dataset (July 2020), only 6.30% of malware samples are unique in the offline-learning setting, and 9.96% are unique in the active-learning setting. This significantly changes the benign-to-malware sample ratio in the deduplicated dataset. Initially, the ratio for this month was 8.81, but it increased to 62.45 for offline learning and 50.65 for active learning deduplication.
We demonstrate in Sections 6 and 7 how the presence of duplicates can undermine the replicability of reported results in Android malware detection, leading to biased evaluations of model performance.
4.2 Model Selection Strategy
@@
A common approach for evaluating machine learning models involves splitting the dataset into training, validation, and test sets. The validation set is essential for model selection and hyperparameter tuning to prevent biased parameter choices [6]. However, previous research on Android malware detection often omits a separate validation set, leading to incomplete evaluations [6]. Our study adopts the same train, validation, and test split as in [15] for offline and active learning settings.
@@
Another common issue is the inadequate tuning of baseline methods [6]. This often results in insufficient hyperparameter tuning for simpler models compared to more complex ones. For example, prior research has shown that appropriately tuned tree-based methods frequently outperform deep neural networks on tabular datasets [24, 39], even though neural networks excel in tasks like computer vision and natural language processing [33, 27, 50, 20].
@@
In Android malware detection, various models have been used in literature, such as support vector machines (SVM), random forests, gradient-boosted decision trees, and deep neural networks [7, 37, 42, 58, 15]. However, these models might not have been adequately calibrated due to the absence of a validation set or insufficient hyperparameter searches. Our analysis focuses on extensive hyperparameter searches for different baseline methods, maintaining the same hyperparameter budget for different models to enable a fair comparison between the models. We detail our procedure in Section 5.3 and show how it can lead to significantly different conclusions about model performance in Section 6.4.
4.3 Accounting for Variance in Performance
@@
Some machine learning models, particularly neural networks, introduce additional stochasticity independent of their hyperparameters. They are sensitive to initial random weight initialization, leading to variations in results based on the training system rather than model parameters [11, 43]. To address this, the deep learning community averages performance across multiple random seeds on the test set and reports the mean performance [12]. This approach captures performance variance, providing a more robust estimate of the model’s capabilities and ensuring reported metrics are not artifacts of a single training instance. However, previous studies on Android malware detection often overlook this variation, resulting in irreproducible outcomes when experiments use different random seeds. We discuss this in detail in Section 6.2.
@@ 4.4 Delayed Evaluation of Models
Splitting the dataset into non-overlapping temporal train, validation, and test sets aids in better model selection. However, this approach often results in the test dataset being delayed from the training dataset. While this is not an issue for classical supervised learning, which assumes an independent and identically distributed setting, it poses challenges for malware detection. The test dataset can be significantly separated from the training dataset by the duration of the validation set, potentially causing concept drift during this period.
For instance, in our study and in [15], there is a six-month gap between the training and test datasets. Using the model trained solely on the training data to evaluate the test set directly might not reflect the practical deployment scenario. Ideally, the model should be trained with more recent data before deployment [6].
To address this characteristic of malware datasets, we propose two different evaluation settings:
  1. 1.
    Merged Training: In this setting, we merge the training and validation datasets after hyperparameter tuning and evaluate the model on the test dataset. We propose that this setting provides a fairer estimate of model performance for offline learning, where the model is already delayed when the test month starts. Note that we do not employ a separate validation set during the merged training phase.
  2. 2.
    Holdout Training: This is the standard setting where the validation dataset is not merged with the training dataset. The test set is separated from the training dataset in this setting, and performance may be underestimated due to concept drift.
We show the effect of these two evaluation settings in Sections 6 and 7.
@@ 5.1 Machine Learning Models
    Support Vector Machine (SVM): SVM is a classification algorithm that finds the hyperplane that best separates classes in the feature space. It has been widely used for Android malware detection, particularly with Drebin features [7, 30].
@@
    Supervised Contrastive Classifier (SCC): Contrastive learning effectively detects and adapts to concept drift in Android malware detection [55, 15]. This approach extracts meaningful representations by distinguishing between similar and dissimilar instance pairs. It assumes that similar instances should be closer in the embedding space while dissimilar ones should be further apart [52, 31].
    We implement a method based on triplet learning, focusing on the main classes—benign and malware—with a triplet loss. In our approach, a sample is considered positive if it belongs to the same class as the anchor and negative if it belongs to a different class. Unlike previous works that treat each malware family as a distinct class, leading to $m+1$ classes where $m$ is the number of malware families, our method only considers the binary case. This is more annotation-friendly, as it does not require separate malware family annotation.
@@
    $\mathcal{L}_{\text{triplet}}=\frac{1}{N}\sum_{i=1}^{N}\max\big{(}$ $0,\|\text{enc}(x_{i}^{a})-\text{enc}(x_{i}^{p})\|_{2}^{2}$
    $-\|\text{enc}(x_{i}^{a})-\text{enc}(x_{i}^{n})\|_{2}^{2}+\text{m}\big{)}$
@@
    Hierarchical Contrastive Classifier (HCC): Chen et al. [15] introduced an enhanced supervised contrastive classifier featuring a hierarchical contrastive loss function. The core idea is to ensure that samples from the same malware family are more similar than those from different families in the feature embedding space. Using an encoder network enc, the embeddings of two benign or two malicious samples from different families should satisfy $\|\text{enc}(x_{1})-\text{enc}(x_{2})\|_{2}\leq m$. For samples from the same malware family, the embeddings should be even more tightly constrained, while embeddings of one malicious and one benign sample should be highly dissimilar, $\|\text{enc}(x_{1})-\text{enc}(x_{2})\|_{2}\geq 2m$.
@@ 5.2 Active Learning Sample Selection Strategy
Active learning focuses on selectively choosing the most informative samples from a large pool of unlabeled data to label, aiming to improve model performance with minimal labeled data. We employ the baseline uncertainty sampling strategy [34] for all models except the HCC. The core idea of uncertainty sampling is to select samples about which the model is least confident. For our models, this involves computing $1-\text{probability}$ (e.g., the softmax output for neural networks) for binary classes and selecting samples with the highest uncertainty.
For HCC, we use pseudo-label sample selection as described in [15]. This method uses model predictions, or pseudo labels, as ground truth to compute the loss function. It calculates the pseudo-loss for contrastive learning by comparing embeddings of test samples against nearby samples in the training dataset. The method then combines the contrastive loss with the binary cross-entropy loss, similar to the one used in training, to identify uncertain samples.
@@ 5.3 Hyperparameter Tuning
We conduct hyperparameter tuning independently for two experimental setups: offline learning and active learning.
In the offline learning setup, each model is trained once on the training dataset, using the validation set to select optimal hyperparameters. Since the model is trained only once, the sample selection strategy (Section 5.2) does not affect this setup. We use random search [10] for hyperparameter tuning, with a fixed budget of 200 searches per model on the validation set. Performance plateaued after 100 iterations for most models. This approach allows for a fair comparison across models and accounts for performance variations due to different hyperparameter search spaces. We perform hyperparameter tuning separately for both the duplicated and deduplicated datasets.
In the active learning setup, we follow the strategy from [15]. The validation data is divided into six-month intervals. We retrain the model using active learning for each interval and evaluate it in the subsequent month. Most hyperparameters, such as batch size for neural networks, remain consistent across the initial training and each retraining phase (a total of five, as no retraining is needed for the last validation month). However, parameters like the number of training epochs during retraining may vary. We use random search with a budget of 100 searches per model and a fixed annotation budget of 50 samples per month, similar to [15], for active learning hyperparameter tuning.
For both setups, we primarily use hyperparameters from previous studies [15], but we have expanded the search space for specific models, such as XGBoost, to improve performance.
Detailed hyperparameters and their influence on validation set performance are provided in the Appendix B.
5.4 Performance Evaluation
After hyperparameter tuning, we evaluate the models on the test set. The active learning setup includes retraining the model each test month and evaluating it the following month. We use the same performance metrics as in [15]: average F1-score, False Positive Rate (FPR), and False Negative Rate (FNR). These metrics differ slightly from the Area Under Time (AUT) metric in [42] but are highly correlated since the interval is consistently one month. Thus, conclusions from both metrics are similar.
Four models, including the random forest and the three neural networks (MLP, SCC, HCC), depend on random seeds for initialization. This reliance makes reproducibility challenging when reimplementing on different platforms, even with identical hyperparameters [60]. To mitigate this issue, we conduct experiments using optimal hyperparameters on the test set with five different seeds, reporting the mean and standard deviation of the metrics. This approach highlights the variability in reproducing model performance across different environments.
While XGBoost can introduce randomness through its sampling methods, we use default parameter values (i.e., 1) to ensure deterministic outcomes for the same model parameters [19]. We confirm deterministic results by running experiments on a different system with the same XGBoost library version.
@@ 5.5 Experimental Settings
We summarize the different experimental settings used in our analysis:
@@
Model Merged Training Holdout Training
Duplicated Deduplicated Duplicated Deduplicated
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
XGBoost 62.9±0.00 0.18±0.00 49.9±0.00 59.9±0.00 0.41±0.00 51.7±0.00 48.0±0.00 0.16±0.00 64.6±0.00 44.9±0.00 0.66±0.00 65.6±0.00
Table 2: Performance comparison of different models on Drebin Dataset for offline learning
Model Merged Training Holdout Training
Duplicated Deduplicated Duplicated Deduplicated
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
RF 64.6±1.02 0.23±0.00 49.6±1.08 61.3±0.92 0.50±0.03 51.7±0.79 45.3±5.2 0.12±0.02 68.5±4.50 46.8±3.89 0.34±0.08 67.0±3.76
XGBoost 74.1±0.00 1.45±0.00 31.7±0.00 68.8±0.00 1.60±0.00 34.7±0.00 70.6±0.00 1.07±0.00 38.5±0.00 66.5±0.00 0.97±0.00 42.7±0.00
Table 3: Performance comparison of different models on APIGraph Dataset for offline learning
@@ 6.1 Malware Detection Performance
We present the results of six models in Table 2 for the Drebin dataset and Table 3 for the APIGraph dataset. These results encompass the various experimental settings discussed previously. Key findings are summarized as follows:
@@
    Impact of Duplicates on Model Performance: Deduplicated datasets generally perform worse than those with duplicates, indicating that duplicates can inflate performance metrics. For instance, in the Drebin dataset’s merged training setting, the F1-score differences for the SVM and SCC models are 17.0 and 17.3, respectively, between duplicate and deduplicated datasets.
@@
    Duplicates Influence Model Preference: Duplicates can skew model comparisons. For example, in the Drebin dataset merge-training setting, SCC outperforms HCC by 6.3 in average F1-score. However, after deduplication, XGBoost outperforms SCC by 2.5 in average F1-score.
@@
    XGBoost’s Optimal Performance: Adequately tuned, XGBoost achieves the highest average F1-score in 3 out of 4 deduplicated datasets and remains competitive in others, suggesting it is a strong baseline model. While SVM performs well on the APIGraph dataset, its performance declines notably on the Drebin dataset after deduplication.
  5. 5.
    Increased Variability Due to Duplicates: Models reliant on random initialization (RF, MLP, SCC, HCC) exhibit greater performance variance on duplicate datasets in most experimental settings. Duplicates not only inflate results but also increase variability, affecting reproducibility. For the APIGraph dataset in the duplicate holdout training setting, HCC’s average F1-score ranges from 57.7 to 70.7, showing a 13-point difference due to random seed variations. We discuss this issue further below.
(a) F1-score on deduplicated data
(b) F1-score on duplicated dataset
(c) FPR on duplicated dataset
(d) FNR on duplicated dataset
Figure 4: Performance in test months on Drebin datasets for Merged Training setting
@@ 6.2 Duplicates and Variance in Performance
Figure 5: Frequency of Unique and top-10 Duplicate malware samples for the month 2015-11 on APIGraph dataset (left) and the month 2021-01 on Drebin dataset (right)
@@
    High Variance Due to Duplicates: As seen in Figure 4(b), duplicates can cause high variance in model performance. Figure 4(d) shows that this effect is primarily due to variance introduced by false negative rates (FNRs). Figure 5(d) illustrates the frequency of unique and top-10 duplicated malware samples for January 2021 on the Drebin dataset. Specifically, 348 out of 385 samples are duplicates of the same input feature, resulting in significant variance for models relying on random initialization, such as neural networks. Two of five MLPs trained with different seeds achieved high F1 scores (93.7 and 94.0), while the others scored much lower (21.4, 24.6, 24.8), indicating high fluctuations due to duplicates. A theoretical classifier correctly identifying all malware samples this month has an FNR of 0. If it fails to detect this single duplicated sample, the FNR rises to 90.4%. Conversely, a classifier detecting no samples has an FNR of 100%, but identifying only the duplicated sample reduces the FNR to 9.6%. This highlights the need to deduplicate datasets for models relying on random initialization to prevent performance disparities caused by "lucky" or "unlucky" initializations.
@@
6.3 Duplicates & Model Selection
The presence of duplicates can also affect hyperparameter tuning and model selection. Figure 6 illustrates this for the XGBoost model, where we adjust the "scale_pos_weight" parameter to address the class imbalance. The positive class weight (malware) is set to 1 in the imbalanced setting. For the balanced setting, it is adjusted to the ratio of benign to malware samples in the training dataset. Balancing the weight improves validation performance for the duplicated dataset in both cases. However, it may reduce performance on the deduplicated APIGraph dataset. This suggests duplicates can cause some hyperparameters to behave differently for the same model.
Figure 7 illustrates how duplicates can influence the model to weigh features differently during training. It compares feature importance between two XGBoost models trained on the deduplicated and the original duplicated Drebin dataset. We extracted the top 50 features based on XGBoost’s feature importance from the deduplicated dataset and evaluated their corresponding importance in the model trained using the duplicated dataset. The figure reveals a disparity in the distribution of feature importance between the two models.

### v2

@@ 4.1.3 Duplicate Statistics in the Datasets
We observe two key patterns regarding duplicates in the validation and test sets. First, the fraction of duplicate samples varies monthly. Second, malware samples tend to contain more duplicates than benign samples, with certain months particularly affected. For instance, in the first test month of the Drebin dataset (July 2020), only 6.30% of malware samples are unique in the offline-learning setting, and 9.96% are unique in the continuous-learning setting.
4.2 Inadequate Model Selection and Tuning
@@
A common approach for evaluating machine learning models involves splitting the dataset into training, validation, and test sets. The validation set is essential for model selection and hyperparameter tuning to prevent biased parameter choices [5]. However, previous research on Android malware detection often omits a separate validation set, leading to incomplete evaluations [5]. Our study adopts the same train, validation, and test split as in [9] for offline and continuous learning settings.
@@
Another common issue is the inadequate tuning of baseline methods [5]. This often results in insufficient hyperparameter tuning for simpler models compared to more complex ones. For example, prior research has shown that appropriately tuned tree-based methods frequently outperform deep neural networks on tabular datasets [38, 39], even though neural networks excel in tasks like computer vision and natural language processing [40, DBLP:conf/nips/VaswaniSPUJGKP17, 41].
@@
In Android malware detection, various models have been used in literature, such as support vector machines (SVM), random forests, gradient-boosted decision trees, and deep neural networks [1, 3, 25, 10, 9]. However, these models might not have been adequately calibrated due to the absence of a validation set or insufficient hyperparameter searches. Our analysis emphasizes extensive hyperparameter searches across different baseline methods, ensuring the same hyperparameter budget is allocated to each model. This approach enables a fair comparison between models and aligns with the established norms in the domain generalization literature in machine learning [42].
4.3 Accounting for Variance from Random Initialization
@@
Some machine learning models, particularly neural networks, exhibit additional stochasticity independent of their hyperparameters. Their sensitivity to initial random weight initialization can lead to variations in results that depend on the training system rather than the model parameters [43, 44]. It is common practice in the deep learning community to evaluate performance across multiple random seeds on the test set and report the average [45] to address this. However, prior studies on Android malware detection often neglect this variation, leading to irreproducible outcomes when experiments are conducted with different random seeds. We explore this issue in detail in Section 6.2.
@@ 4.4 Delayed Evaluation of Models
Splitting the dataset into non-overlapping temporal train, validation, and test sets enables consistent model selection but introduces a delay between training and test data. While acceptable in standard supervised learning, this temporal gap poses challenges for malware detection, where data distributions evolve. As observed in our study and in [9], a six-month delay between training and test periods can lead to unrealistic evaluations, since deployed models would typically be retrained with more recent data [5].
To mitigate this, we merge the training and validation datasets after hyperparameter tuning and evaluate on the test set. This Merged Training setup offers a fairer estimate of offline performance compared to the Holdout Training setup, where only the training data are used. Such merging is standard in temporal domain generalization [46], and the impact of both setups is analyzed in Sections 6 and 7.
@@ 5.1 Machine Learning Models
    Support Vector Machine (SVM): SVM [48] is a classification algorithm that finds the hyperplane that best separates classes in the feature space. It has been widely used for Android malware detection, particularly with Drebin features [1, 24].
@@
    Supervised Contrastive Classifier (SCC): Contrastive learning has proven effective for detecting and adapting to concept drift in Android malware detection [14, 9]. It learns representations by bringing similar instances closer and pushing dissimilar ones apart [51, 52]. Our implementation uses a triplet loss over the two main classes, benign and malware, where positives share the same label as the anchor and negatives differ. Unlike prior works that model each malware family as a separate class, our binary setup is more annotation efficient and better suited for real world detection.
@@
    $\mathcal{L}_{\text{triplet}}$ $=\frac{1}{N}\sum_{i=1}^{N}\max\Bigl(0,\;\|\mathrm{enc}(x_{i}^{a})-\mathrm{enc}(x_{i}^{p})\|_{2}^{2}$
    $\quad\mathllap{-}\;\|\mathrm{enc}(x_{i}^{a})-\mathrm{enc}(x_{i}^{n})\|_{2}^{2}+m\Bigr)$
@@
    Hierarchical Contrastive Classifier (HCC): Following [9], HCC extends supervised contrastive learning with a hierarchical loss that enforces stronger similarity among samples from the same malware family than those from different families. Using an encoder $\mathrm{enc}$, embeddings of benign or cross-family malware pairs satisfy $\|\mathrm{enc}(x_{1})-\mathrm{enc}(x_{2})\|_{2}\leq m$, while benign–malicious pairs are pushed farther apart with $\|\mathrm{enc}(x_{1})-\mathrm{enc}(x_{2})\|_{2}\geq 2m$.
@@ 5.2 Active Learning Sample Selection Strategy
Active learning seeks to improve model performance with minimal labeled data by selectively querying the most informative samples from a large pool of unlabeled data [53]. For all models except HCC, we employ the standard uncertainty sampling strategy [54], which selects samples where the model is least confident, computed as $1-\text{probability}$ (e.g., the softmax output for binary classifiers). For HCC, we follow the pseudo-label selection method of [9], where pseudo-labels act as ground truth for calculating a contrastive pseudo-loss between test and nearby training embeddings. This loss is combined with binary cross-entropy to identify samples with the highest uncertainty.
TABLE II: Performance comparison of different models on the Drebin dataset for offline learning. The results represent the mean and standard deviation over five runs with different random seeds. The best-performing model is underlined.
Model Merged Training Holdout Training
Duplicated Deduplicated Duplicated Deduplicated
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
XGBoost 63.7±2.45 0.18±0.01 48.9±2.80 59.5±0.46 0.39±0.02 52.1±0.59 45.2±1.92 0.24±0.04 67.3±2.01 47.3±0.78 0.73±0.02 63.4±0.66
TABLE III: Performance comparison of different models on the APIGraph Dataset for offline learning
Model Merged Training Holdout Training
Duplicated Deduplicated Duplicated Deduplicated
F1 FPR FNR F1 FPR FNR F1 FPR FNR F1 FPR FNR
RF 64.6±1.02 0.23±0.00 49.6±1.08 61.3±0.92 0.50±0.03 51.7±0.79 45.3±5.22 0.12±0.02 68.5±4.50 46.8±3.89 0.34±0.08 67.0±3.76
XGBoost 74.6±0.47 1.21±0.09 32.5±0.49 69.0±0.45 1.61±0.03 34.3±0.57 69.5±1.40 1.03±0.06 40.2±1.65 65.2±0.62 1.33±0.05 41.6±0.88
@@ 5.3 Hyperparameter Tuning
We perform hyperparameter tuning independently for two experimental setups: offline learning and continuous active learning. In the offline setup, each model is trained once on the training dataset, using the validation set to select optimal hyperparameters. We apply random search [55] with a budget of 200 trials per model, ensuring fair comparison across models and accounting for differences in search spaces, and repeat tuning separately for duplicated and deduplicated datasets. In the continuous active learning setup, following [9], the validation data is divided into six-month intervals; the model is retrained with active learning at each interval and evaluated on the subsequent month. Most hyperparameters remain fixed across retraining phases, though parameters such as training epochs may vary. Random search with 100 trials per model and an annotation budget of 50 samples per month is used for tuning. Full hyperparameter details are provided in Appendix A.
5.4 Evaluation Metric
After hyperparameter tuning, we evaluate the models on the test set. The continuous active learning setup includes retraining the model each test month and evaluating it the following month. We use the same performance metrics as in [9]: average F1-score, False Positive Rate (FPR), and False Negative Rate (FNR). We evaluate all models on the test set using optimal hyperparameters and five different random seeds, reporting the mean and standard deviation of the metrics. It is worth noting that SVM produces deterministic results across seeds, as it uses the same training dataset without any stochastic elements.
@@ 5.5 Experimental Settings
We summarize the different experimental settings below:
@@
(a) F1-score on deduplicated data
(b) F1-score on duplicated dataset
(c) FNR on duplicated dataset
Figure 4: Performance in test months on the Drebin datasets for Merged Training setting
@@ 6.1 Malware Detection Performance
We present the results of six models in Table II for the Drebin dataset and Table III for the APIGraph dataset. These results encompass the various experimental settings discussed previously. Key findings are summarized as follows:
@@
    Impact of duplicates on model performance: Deduplicated datasets generally perform worse than those with duplicates, indicating that duplicates may inflate performance metrics. For instance, in the Drebin dataset’s merged training setting, the F1-score differences for the SVM and SCC models are 17.0 and 17.3, respectively, between duplicate and deduplicated datasets.
@@
    Duplicates influence model preference: Duplicates can skew model comparisons. For example, in the Drebin dataset merge-training setting, SCC outperforms XGBoost by 11 in average F1-score. However, after deduplication, XGBoost outperforms SCC by 2.1 in average F1-score.
@@
    XGBoost as a strong baseline on deduplicated datasets: With proper tuning, XGBoost achieves the highest average F1-score on the deduplicated Drebin dataset and remains competitive on APIGraph, highlighting its effectiveness as a baseline model. In contrast, while SVM performs well on APIGraph, its performance drops significantly on the Drebin dataset after deduplication.
Figure 5: Frequency of unique and duplicate malware samples. Left: APIGraph in November 2015 with 190 of 488 samples sharing identical features. Right: Drebin in January 2021, with 348 of 385 samples duplicates of the same input.
@@ 6.2 Duplicates and Variance in Performance
    High variance due to duplicates: As seen in Figure 4(b), duplicates can cause high variance in model performance. Figure 4(c) shows that this effect is primarily due to variance introduced by false negative rates (FNRs). Figure 5(right) illustrates the frequency of unique, and top-10 duplicated malware samples for January 2021 on the Drebin dataset. Specifically, 348 out of 385 samples are duplicates of the same input feature, resulting in significant variance for models relying on random initialization, such as neural networks. Two of five MLPs trained with different seeds achieved high F1 scores (93.7 and 94.0), while the others scored much lower (21.4, 24.6, 24.8), indicating high fluctuations due to duplicates. A theoretical classifier correctly identifying all malware samples this month has an FNR of 0. If it fails to detect this single duplicated sample, the FNR rises to 90.4%. Conversely, a classifier detecting no samples has an FNR of 100%, but identifying only the duplicated sample reduces the FNR to 9.6%. This highlights the need to deduplicate datasets for models relying on random initialization to prevent performance disparities caused by different initializations.
@@
6.3 Effective Model Selection for Robust Baselines
Our offline learning experiments show that XGBoost consistently outperforms other models across settings, differing from the results in [9] primarily due to an expanded hyperparameter search space. We included additional parameters and wider ranges (see Appendix A), with the number of boosting rounds being particularly important. While [9] tested only 10–100 rounds, we found that up to 400 rounds significantly improved performance, especially on the APIGraph dataset. This highlights the importance of thorough hyperparameter tuning for establishing strong baselines [5]. Figure 6 shows the impact of boosting rounds on validation performance across datasets.